OpenStudy (anonymous):

Look at the formula for finding distance. d = rt A. t = d --- r B. t = d - r C. t = r --- d D. t = dr How do you do this?

sam (.sam.):

assume r=velocity d=rt

sam (.sam.):

standard formula for velocity is \[v=\frac{d}{t}\] Multiply both sides by 't' d=vt

OpenStudy (anonymous):

I don't understand, sorry. This is very confusing.

sam (.sam.):

which grade you in?

sam (.sam.):

you mean you want to find "t"?

OpenStudy (anonymous):

how do you multiply a letter by a letter is what i don't understand -.-

sam (.sam.):

Just divide both sides by "t" \[d = rt\] \[\frac{d}{r}=\frac{rt}{r}\] \[\frac{d}{r}=\frac{\cancel rt}{\cancel r}\] \[\frac{d}{r}=t\]

sam (.sam.):

I mean divide both sides by "r"

sam (.sam.):

So its A

OpenStudy (anonymous):

@michaelJordan you are not multiplying by a letter.they are symbols that is used to represent a variable value.its a number afterall.

OpenStudy (anonymous):

k